OpenStudy (maddy1251):

Complete the following table without using your calculator: theta = 30 degrees sin theta =? cos theta = ? tan theta = ?

OpenStudy (anonymous):

 take the square root of each. also tan=sin/cos

OpenStudy (anonymous):

In a 30 degree right triangle, the side opposite the 30 degree angle is 1/2 the length of the hypotenuse. Let the hypotenuse length be one, 1 . Then the sine of 30 deg is (1/2)/1 or 1/2 The cosine is the length of the third side or \[\sqrt{1^2-\left(\frac{1}{2}\right)^2}=\frac{\sqrt{3}}{2} \]The Tangent is the ratio of the sine to the cosine.